# Exploring Cultural Variations in Moral Judgments with Large Language Models

### v1

@@ Abstract
Large Language Models (LLMs) have shown strong performance across many tasks, but their ability to capture culturally diverse moral values remains unclear. In this paper, we examine whether LLMs can mirror variations in moral attitudes reported by two major cross-cultural surveys: the World Values Survey and the PEW Research Center’s Global Attitudes Survey. We compare smaller, monolingual, and multilingual models (GPT-2, OPT, BLOOMZ, and Qwen) with more recent instruction-tuned models (GPT-4o, GPT-4o-mini, Gemma-2-9b-it, and Llama-3.3-70B-Instruct). Using log-probability-based moral justifiability scores, we correlate each model’s outputs with survey data covering a broad set of ethical topics. Our results show that many earlier or smaller models often produce near-zero or negative correlations with human judgments. In contrast, advanced instruction-tuned models (including GPT-4o and GPT-4o-mini) achieve substantially higher positive correlations, suggesting they better reflect real-world moral attitudes. While scaling up model size and using instruction tuning can improve alignment with cross-cultural moral norms, challenges remain for certain topics and regions. We discuss these findings in relation to bias analysis, training data diversity, and strategies for improving the cultural sensitivity of LLMs.
Exploring Cultural Variations in Moral Judgments
with Large Language Models
Hadi Mohammadi, Efthymia Papadopoulou, Yasmeen F.S.S. Meijer, and Ayoub Bagheri Department of Methodology and Statistics, Utrecht University, Utrecht, The Netherlands h.mohammadi@uu.nl, evi.papado98@gmail.com, mijntje.meijer@live.nl ,and a.bagheri@uu.nl
@@ 1 Introduction
Over the past few years, LLMs have gained prominence in both academic and public discussions (Bender et al., 2021). Advances in model performance have made LLMs appealing for diverse applications, such as social media content moderation, chatbots, content creation, real-time translation, search engines, recommendation systems, and automated decision-making. While modern LLMs (e.g., GPT-4) show strong performance, a critical concern is how these models may inherit biases, including gender, racial, or cultural biases, from their training data. LLMs can easily absorb such biases because they learn from large-scale text corpora containing entrenched stereotypes (Stańczak and Augenstein, 2021; Karpouzis, 2024).
@@
These biases raise concerns about fairness, particularly in contexts requiring moral judgments. If an LLM is trained mostly on data that negatively or inaccurately portrays certain cultural groups, it may repeat that bias in its responses. As these models become more widespread and globally deployed, the risk of perpetuating cultural biases grows, especially when moral perspectives are different from common rules or what surveys usually show. In fact, recent research shows that current LLMs often exhibit a default Western-centric bias (Adilazuarda et al., 2024), underscoring the need to evaluate their cross-cultural validity
@@
It is crucial to see whether LLMs accurately mirror the moral judgments observed across diverse cultures. Despite its importance, this issue has received limited attention (Arora et al., 2023; Liu et al., 2024). Our study investigates whether both monolingual and multilingual Pre-trained Language Models (PLMs) can capture nuanced cultural norms. These norms include subtle ethical differences across regions, for example, the acceptance of alcohol consumption or differing attitudes on topics like abortion. Although recent research suggests that multilingual PLMs might capture broader cultural nuances, they often fall short of reflecting the moral subtleties present in less dominant cultural groups (Hämmerl et al., 2022; Papadopoulou et al., 2024).
@@
We examine this question using two well-known cross-cultural datasets: the World Values Survey (WVS) (Inglehart et al., 2014; Haerpfer et al., 2022), and the PEW Research Center’s Global Attitudes Survey, which includes a module on moral issues across many countries (Pew Research Center, 2023). These surveys offer a detailed view of moral and cultural norms globally, serving as a benchmark for comparing LLMs outputs against actual human responses. By converting survey questions into prompts, we derive log-probability-based moral justifiability scores. We then compare these scores with survey-based consensus on various ethical issues (e.g., drinking alcohol, sex before marriage, abortion, homosexuality), allowing us to see how closely different model types and training approaches align with cultural norms. Evaluating how effectively LLMs represent cultural values has both scholarly and practical significance. If a model systematically misrepresents or overlooks certain moral perspectives, it may reinforce stereotypes or lead to biased outcomes. On the other hand, more culturally aware models can highlight both shared values and nuanced disagreements, potentially contributing to more balanced dialogue. By comparing model outputs to reliable survey data, we identify areas where LLMs align with human values and highlight gaps in capturing diverse moral perspectives.
@@
Our contributions are threefold: (1) We introduce a structured probing framework that leverages carefully designed prompts, contrasting moral statements, and log-probability-based scoring to assess how LLMs assign justifiability values to morally complex scenarios across cultures. (2) We empirically analyze the alignment between LLM-derived moral scores and human survey responses using correlation and clustering, highlighting where models reflect or deviate from real-world moral judgments. (3) We extend our evaluation to state-of-the-art instruction-tuned and large-scale models, examining whether instruction tuning and scaling enhance alignment with cross-cultural moral norms. By identifying key strengths, weaknesses, and factors influencing model-human agreement, our work contributes to improving training data strategies, mitigating biases, and fostering the development of culturally aware language models.
@@ 2 Literature review
LLMs inherit biases embedded in their training data, and these biases can be amplified upon large-scale deployment. Because the underlying corpora often reflect entrenched social hierarchies, models run the risk of reproducing or even intensifying unfair patterns. Recent work has underscored this from multiple perspectives, a 2025 study introduced a unified framework for transparency, fairness, and privacy in AI pipelines (Radanliev, 2025), while an interdisciplinary survey emphasized the importance of diversity, equity, and inclusion (DEI) as prerequisites for trustworthy AI (Cachat-Rosset and Klarsfeld, 2023a). Taken together with earlier warnings about opaque language-model behaviors (Bender et al., 2021), these findings illustrate the need for technical innovation to proceed hand-in-hand with social safeguards. In addition to high-level ethical governance, researchers are exploring concrete mitigation strategies. For example, LLM data augmentation has improved intent-classification accuracy without sacrificing fairness, provided that the augmentation is carefully curated (Benayas et al., 2024). Complementary work on adapter tuning for non-English LLMs shows that modest architectural modifications can substantially boost performance in culturally diverse benchmarks, thereby supporting more inclusive NLP systems (Zhou et al., 2024).
@@
Moral judgments themselves—evaluations of actions, intentions, or individuals as acceptable or objectionable—can differ widely by culture, shaped by religious traditions, social norms, and historical contexts (Haidt, 2001; Shweder et al., 1997). Understanding how such pluralistic values are (or are not) embedded in contemporary LLMs remains a pressing research concern. As noted by Graham et al. (2016), Western, Educated, Industrialized, Rich, and Democratic (W.E.I.R.D.) societies emphasize individual rights and autonomy, while non-W.E.I.R.D. societies often stress communal responsibilities and spiritual considerations. Consequently, people in W.E.I.R.D. cultures may view personal choices like sexual behavior as an individual right, while those in non-W.E.I.R.D. cultures consider them a collective moral concern. Although many moral values overlap across cultures, there are also areas of genuine divergence, often referred to as moral value pluralism (Johnson et al., 2022; Benkler et al., 2023). However, Kharchenko et al. (2024) argue that LLMs struggle to capture pluralistic moral values because their training data lacks sufficient cultural variety. Likewise, Du et al. (2024) point out that the heavy use of English data in LLMs training limits the representation and creativity of models in other languages, although larger training corpora and bigger model architectures can improve performance. Arora et al. (2023) suggest that multilingual LLMs could learn cultural values by incorporating multilingual data in their training. Yet, the limited diversity within multilingual corpora can still cause these models to perform inconsistently across languages and cultural contexts. Benkler et al. (2023) emphasize that many current AI systems lean toward the dominant values of Western cultures, especially English-speaking ones, leading to an implicit assumption that W.E.I.R.D. values are universal.
@@
During training, LLMs use word embeddings to learn semantic and syntactic relationships based on how frequently words co-occur. These embeddings can encode the same social biases found in the training data (Nemani et al., 2024; Mohammadi et al., 2025). This association-based learning can produce biased outputs that influence the model’s fairness and reliability. For instance,  Johnson et al. (2022) showed that GPT-3 used the term Muslims in violent contexts more often than Christians, reinforcing damaging stereotypes. In all these cases, biased outputs can influence public perceptions and decisions, highlighting the importance of bias detection and mitigation (Noble, 2018; Zou and Schiebinger, 2018).
@@
Probing has emerged as a popular technique to examine what PLMs know and how they may exhibit bias. Ousidhoum et al. (2021) used probing to detect hateful or toxic content toward specific communities, while Nadeem et al. (2021) used context-based association tests to investigate stereotypes. Arora et al. (2023) adapted cross-cultural survey questions into prompts to test multilingual PLMs in 13 languages, discovering that these models often failed to match the moral values embedded in their training languages. Although there are multiple probing approaches, from cloze-style tasks to pseudo-log-likelihood scoring (Nadeem et al., 2021; Salazar et al., 2019), each has limitations. A simpler method directly computes the probability of specific tokens, following the original transformer design (Vaswani et al., 2017).
@@
Research on AI ethics underscores the need for models that respect cultural distinctions and support equitable treatment (Zowghi and da Rimini, 2023; Cachat-Rosset and Klarsfeld, 2023b; Karpouzis, 2024; Meijer et al., 2024). Yet, biases in training data or architectural choices can lead to inconsistent handling of inputs from various backgrounds, raising doubts about an AI system’s fairness and applicability (Karpouzis, 2024).While studies like Arora et al. (2023) and Benkler et al. (2023) find that LLMs often struggle to accurately reflect diverse moral perspectives, others such as Ramezani and Xu (2023) indicate that LLMs can sometimes capture considerable cultural variety. Similarly, Cao et al. (2023) showed that ChatGPT aligns strongly with American cultural norms while adapting less effectively to others, reinforcing concerns of Western-centric bias in LLM outputs. This discrepancy highlights the need for more research on how LLMs learn and represent moral values in different cultural settings. Even though LLMs can inherit some cultural biases, the extent of their cross-cultural fidelity remains an open question (Caliskan et al., 2017).
@@ 3.1 Data
To evaluate cross-cultural moral attitudes, we use two datasets: World Values Survey (WVS) Wave 7 and the PEW Research Center Global Attitudes Survey 2013. Each dataset’s moral questions are labeled with topic codes. See Table 4 in Appendix A for a full reference.
World Values Survey Wave 7
@@
The WVS conducted from 2017 to 2020111https://www.worldvaluessurvey.org/WVSDocumentationWV7.jsp, which covers respondents from 55 countries (Inglehart et al., 2014; Haerpfer et al., 2022). We use the section of the survey dealing with Ethical Values and Norms. In this section, participants were asked to rate the justifiability of 19 different behaviors or issues with moral connotations. These include topics such as divorce, euthanasia, political violence, cheating on taxes, and others. We performed preprocessing by filtering the dataset to retain only the responses to the 19 moral questions (Q177 to Q195) and the country code for each respondent.
@@
Each response is an integer from 1 to 10. We then mapped the country codes to country names (using the provided codebook) so that each respondent entry includes their country and their answers to the moral questions. Next, we handled missing or non-response values. Entries coded as $-1$, $-2$, $-4$, or $-5$ (i.e., Don’t know, No answer, Not asked, and Missing) were set to 00, so they would not distort later calculations. We then grouped the data by country and averaged the responses for each moral statement. This yields a country-level average moral approval score for each of the 19 issues. Because different countries may use the $1$-–$10$ scale differently (culturally, some may avoid extreme ratings, etc.), and to facilitate comparison with the second dataset, we normalized these country mean scores to a range of $[-1,1]$, with $-1$ denoting never justifiable and $+1$ denoting always justifiable.
After these steps, the WVS data provides, for each country and each moral topic, a score between -1 and 1 representing how acceptable that behavior is on average according to that country’s respondents. Higher scores mean the society tends to view the behavior as more acceptable or justifiable, whereas lower scores mean it is seen as less acceptable or not justifiable. We treat these normalized country-by-topic scores as the empirical ground truth of moral attitudes.
Figure 1: Spread of responses (country mean scores) across the moral topics for the WVS Wave 7 dataset.
Figure 1 shows the spread of responses across different moral topics and countries. In other words, for each moral topic, how varied are the country scores? Some topics might have very similar scores in every culture (indicating global agreement), while others show a wide range (indicating high cross-cultural controversy).
PEW Global Attitudes Survey 2013
@@
The PEW collected responses on moral issues from 39 countries, with about 100 respondents per country for the relevant questions [^2] https://www.pewresearch.org/dataset/spring-2013-survey-data/. Unlike WVS, which used a 10-point scale, the PEW survey questions were simpler: for each issue, respondents were asked whether the behavior is morally acceptable, morally unacceptable, or not a moral issue.
@@
From the PEW dataset, we extracted the questions corresponding to those eight moral topics (Q84A to Q84H). We again retained only the country identifier and these responses for our purposes. We coded the responses in a numeric way to be analogous to the WVS scale: for each question, we assigned a value of $+1$ to morally acceptable, $-1$ to morally unacceptable, and 00 to not a moral issue and all non-responses (including Depends on situation, Refused, and Don’t know). As with WVS, we grouped responses by country, averaged them for each topic, and normalized the averages to $[-1,1]$. Figure 2 shows the normalized PEW values across the eight moral questions. The comparison of normalized scores for WVS and PEW by country is also presented in Appendix B, Figure 8.
Figure 2: Spread of responses across the moral topics and countries for the PEW 2013 dataset.

### v2

@@ Abstract
Large Language Models (LLMs) have shown strong performance across many tasks, but their ability to capture culturally diverse moral values remains unclear. In this paper, we examine whether LLMs mirror variations in moral attitudes reported by the World Values Survey (WVS) and the Pew Research Center’s Global Attitudes Survey (PEW). We compare smaller monolingual and multilingual models (GPT-2, OPT, BLOOMZ, and Qwen) with recent instruction-tuned models (GPT-4o, GPT-4o-mini, Gemma-2-9b-it, and Llama-3.3-70B-Instruct). Using log-probability-based moral justifiability scores, we correlate each model’s outputs with survey data covering a broad set of ethical topics. Our results show that many earlier or smaller models often produce near-zero or negative correlations with human judgments. In contrast, advanced instruction-tuned models achieve substantially higher positive correlations, suggesting they better reflect real-world moral attitudes. We provide a detailed regional analysis revealing that models align better with Western, Educated, Industrialized, Rich, and Democratic (W.E.I.R.D.) nations than with other regions. While scaling model size and using instruction tuning improves alignment with cross-cultural moral norms, challenges remain for certain topics and regions. We discuss these findings in relation to bias analysis, training data diversity, information retrieval implications, and strategies for improving the cultural sensitivity of LLMs.
Keywords: Large Language Models, Moral judgments, Cultural variation, Cross-cultural analysis, Bias detection
@@ 1 Introduction
Over the past few years, Large Language Models (LLMs) have gained prominence in both academic and public discussions (?). Consider this example: when asked about the moral acceptability of divorce, an LLM might predict similar attitudes for Sweden and Saudi Arabia—yet survey data reveals nearly opposite positions on a normalized scale. Such blind spots matter as LLMs increasingly power content moderation, search engines, and decision-support systems globally. Advances in model performance have made LLMs appealing for diverse applications, such as social media content moderation, chatbots, content creation, real-time translation, search engines, recommendation systems, and automated decision-making. While modern LLMs (e.g., GPT-4) show strong performance, a critical concern is how these models may inherit biases, including gender, racial, or cultural biases, from their training data. LLMs can easily absorb such biases because they learn from large-scale text corpora containing entrenched stereotypes (?, ?). Recent work further examines whether LLMs capture cross-cultural moral judgments, reporting partial alignment alongside systematic blind spots across topics and regions (?). Our study complements these findings by correlating model-derived justifiability scores with the World Values Survey (WVS) and the Pew Research Center’s Global Attitudes Survey (PEW) across a broader mix of model families and elicitation settings.
@@
These biases raise concerns about fairness, particularly in contexts requiring moral judgments. If an LLM is trained mostly on data that negatively or inaccurately portrays certain cultural groups, it may repeat that bias in its responses. As these models become more widespread and globally deployed, the risk of perpetuating cultural biases increases, especially when moral perspectives diverge from common norms or survey results. Recent research indicates that current LLMs often exhibit a Western-centric bias (?), underscoring the need to evaluate their cross-cultural validity. A large-scale evaluation across 107 countries found that cultural prompting reduces bias for approximately 71–81% of countries but fails for the remaining 19–29% (?).
@@
It is crucial to determine whether LLMs accurately mirror the moral judgments observed across diverse cultures. Despite its importance, this issue has received limited attention (?, ?). Our study investigates whether both monolingual and multilingual Pre-trained Language Models (PLMs) can capture nuanced cultural norms. These norms include subtle ethical differences across regions, for example, the acceptance of alcohol consumption or differing attitudes on topics like abortion. Although recent research suggests that multilingual PLMs might capture broader cultural nuances, they often fall short of reflecting the moral subtleties present in less dominant cultural groups (?, ?).
@@
We examine this question using two well-known cross-cultural datasets: the WVS (?, ?), and the PEW, which includes a module on moral issues across many countries (?). These surveys offer a detailed view of global moral and cultural norms, serving as a benchmark for comparing LLM outputs against human responses. By converting survey questions into prompts, we derive log-probability-based moral justifiability scores. We then compare these scores with survey-based consensus on various ethical issues (e.g., drinking alcohol, sex before marriage, abortion, homosexuality), allowing us to see how closely different model types and training approaches align with cultural norms.
Evaluating how effectively LLMs represent cultural values has both scholarly and practical significance. If a model systematically misrepresents certain moral perspectives, it may reinforce stereotypes or lead to biased outcomes. This has direct implications for Information Retrieval (IR) systems: as LLMs are increasingly integrated into search engines, recommendation systems, and content moderation pipelines, their cultural biases can affect information access. A model that underrepresents moral perspectives from certain regions may systematically disadvantage users from those regions in personalized search, content filtering, or cross-cultural information needs. Conversely, culturally aware models can highlight shared values and nuanced disagreements, potentially contributing to more balanced dialogue.
@@
Our contributions are fourfold: (1) We introduce a structured probing framework that leverages carefully designed prompts, contrasting moral statements, and log-probability-based scoring to assess how LLMs assign justifiability values to morally complex scenarios across cultures. (2) We empirically analyze the alignment between LLM-derived moral scores and human survey responses using correlation and clustering, highlighting where models reflect or deviate from real-world moral judgments. (3) We provide a detailed regional analysis comparing model performance across W.E.I.R.D. versus non-W.E.I.R.D. nations and different geographical regions. (4) We extend our evaluation to state-of-the-art instruction-tuned and large-scale models, examining whether instruction tuning and scaling enhance alignment with cross-cultural moral norms.
@@ 2 Literature Review
LLMs inherit biases embedded in their training data, and these biases can be amplified upon large-scale deployment. Because the underlying corpora often reflect entrenched social hierarchies, models run the risk of reproducing or even intensifying unfair patterns. Recent work has underscored this from multiple perspectives. For example, a 2025 study introduced a unified framework for transparency, fairness, and privacy in Artificial Intelligence (AI) pipelines (?), while an interdisciplinary survey emphasized the importance of diversity, equity, and inclusion as prerequisites for trustworthy AI (?). Taken together with earlier warnings about opaque language-model behaviors (?), these findings illustrate the need for technical innovation alongside social safeguards. In addition to high-level ethical governance, explainability-oriented methods can support bias analysis and mitigation at the text level (?). For instance, explanation-guided token replacement has been shown to steer LLM outputs away from problematic attributions while preserving utility (?). Beyond dataset and architectural factors, the reliability of LLM-produced judgments and annotations is itself uneven across demographic slices (?), reinforcing the need for careful evaluation protocols alongside fairness safeguards.
@@
Moral judgments—evaluations of actions, intentions, or individuals as acceptable or objectionable—differ widely by culture, shaped by religious traditions, social norms, and historical contexts (?, ?). Understanding how pluralistic values are embedded in contemporary LLMs remains a pressing research concern. As noted by ?, W.E.I.R.D. societies emphasize individual rights and autonomy, while non-W.E.I.R.D. societies often stress communal responsibilities and spiritual considerations. Consequently, people in W.E.I.R.D. cultures may view personal choices like sexual behavior as an individual right, while those in non-W.E.I.R.D. cultures consider them a collective moral concern.
Although many moral values overlap across cultures, there are also areas of genuine divergence, often referred to as moral value pluralism (?, ?). However, ? argue that LLMs struggle to capture pluralistic moral values because their training data lacks sufficient cultural variety. Likewise, ? point out that the heavy use of English data in LLMs training limits the representation and creativity of models in other languages, although larger training corpora and bigger model architectures can improve performance. Recent work by ? demonstrates that ethical reasoning and moral value alignment in LLMs depend significantly on the language used in prompts, with models showing different moral positions when queried in different languages. ? suggest that multilingual LLMs could learn cultural values by incorporating multilingual data in their training. Yet, the limited diversity within multilingual corpora can still cause these models to perform inconsistently across languages and cultural contexts. ? emphasize that many current AI systems lean toward the dominant values of Western cultures, especially English-speaking ones, leading to an implicit assumption that W.E.I.R.D. values are universal.
Recent work has explored alternative theoretical frameworks for analyzing moral alignment in LLMs. ? applied Moral Foundations Theory to examine how moral biases vary across different prompting contexts, while ? developed utilitarian dilemma-based benchmarks that complement survey-based approaches. ? found that Chinese LLMs exhibit collectivist values compared to Western models’ individualistic tendencies, demonstrating that model origin significantly influences moral outputs. Most recently, ? introduced WorldValuesBench, a large-scale benchmark specifically designed to evaluate multi-cultural value awareness in LLMs using WVS data, providing systematic evaluation across diverse cultural contexts.
@@
During training, LLMs develop associations between concepts based on co-occurrence patterns in text. These learned associations can encode the same social biases found in the training data (?, ?). This association-based learning can produce biased outputs that influence the model’s fairness and reliability. For instance, ? showed that GPT-3 used the term Muslims in violent contexts more often than Christians, reinforcing damaging stereotypes. In all these cases, biased outputs can influence public perceptions and decisions, highlighting the importance of bias detection and mitigation (?, ?).
@@
Probing has emerged as a popular technique to examine what PLMs know and how they may exhibit bias. ? used probing to detect hateful or toxic content toward specific communities, while ? used context-based association tests to investigate stereotypes. ? adapted cross-cultural survey questions into prompts to test multilingual PLMs in 13 languages, discovering that these models often failed to match the moral values embedded in their training languages. Although there are multiple probing approaches, from cloze-style tasks to pseudo-log-likelihood scoring (?, ?), each has limitations. A simpler method directly computes the probability of specific tokens, following the original transformer design (?).
2.1 Relationship to Prior Work
Our work builds upon the foundational framework introduced by ?, who first systematically evaluated whether LLMs contain knowledge about moral norms across cultures using WVS data. We extend their approach in several important ways, as summarized in Table 1.
Table 1: Comparison of our work with ?.
Aspect Ramezani & Xu (2023) Our Work
Models evaluated 5 models 20 models across 9 families
Datasets WVS only WVS + PEW (cross-validation)
Analysis type Binary knowledge assessment Correlation + clustering + error analysis
Instruction-tuned models Not evaluated GPT-4o, Gemma-2, Falcon-40B-Inst
Topic difficulty analysis Not included Easy vs. hard topics identified
Regional analysis Limited W.E.I.R.D. vs. non-W.E.I.R.D. breakdown
@@ 2.1 Relationship to Prior Work
Research on AI ethics underscores the need for models that respect cultural distinctions and support equitable treatment (?, ?, ?, ?). Yet, biases in training data or architectural choices can lead to inconsistent handling of inputs from various backgrounds, raising doubts about an AI system’s fairness and applicability (?). While studies like ? and ? find that LLMs often struggle to accurately reflect diverse moral perspectives, others such as ? indicate that LLMs can sometimes capture considerable cultural variety. Similarly, ? showed that ChatGPT aligns strongly with American cultural norms while adapting less effectively to others, reinforcing concerns of Western-centric bias in LLM outputs.
Having established this theoretical landscape and identified key gaps in prior work, we now describe our methodology for systematically evaluating cultural moral alignment across 20 models and 63 countries.
@@ 3.1 Data
To evaluate cross-cultural moral attitudes, we use two datasets: WVS Wave 7 and the PEW Global Attitudes Survey 2013. Table 2 provides summary statistics for both datasets.
Table 2: Dataset summary statistics.
Dataset Countries Topics Scale Description
WVS Wave 7 55 19 1–10 (numeric) Conducted 2017–2020; covers ethical values and norms
PEW 2013 39 8 Categorical Morally acceptable, unacceptable, or not a moral issue
World Values Survey Wave 7.
@@
The WVS, conducted from 2017 to 2020, covers respondents from 55 countries (?, ?). We use the section dealing with Ethical Values and Norms, where participants rated the justifiability of 19 different behaviors or issues with moral connotations. These include topics such as divorce, euthanasia, political violence, cheating on taxes, and others. We performed preprocessing by filtering the dataset to retain only the responses to the 19 moral questions (Q177 to Q195) and the country code for each respondent.
@@
Each response is an integer from 1 to 10. We mapped the country codes to country names using the provided codebook. For missing or non-response values (codes 1-1, 2-2, 4-4, or 5-5 representing Don’t know, No answer, Not asked, and Missing), we excluded these responses from calculations rather than coding them as zero, to avoid artificially biasing the mean estimates. We then grouped the data by country and averaged the responses for each moral statement. This yields a country-level average moral approval score for each of the 19 issues. To facilitate comparison with the second dataset, we normalized these country mean scores to a range of [1,1][-1,1], with 1-1 denoting never justifiable and +1+1 denoting always justifiable.
We acknowledge that this min-max normalization does not fully address differences in how various cultures may use rating scales (e.g., some cultures may avoid extreme ratings). This remains a limitation of our approach.
PEW Global Attitudes Survey 2013.
@@
The PEW collected responses on moral issues from 39 countries, with approximately 100 respondents per country for the relevant questions. Unlike WVS, which used a 10-point scale, the PEW survey questions were simpler: for each issue, respondents were asked whether the behavior is morally acceptable, morally unacceptable, or not a moral issue.
@@
From the PEW dataset, we extracted the questions corresponding to eight moral topics (Q84A to Q84H). We coded the responses as follows: +1+1 for morally acceptable, 1-1 for morally unacceptable, and excluded non-responses (including Depends on situation, Refused, and Don’t know) from calculations. As with WVS, we grouped responses by country, averaged them for each topic, and normalized the averages to [1,1][-1,1].